# Collective optical properties of moiré excitons

### v1

@@ Abstract
Berry curvature. — Topological features can also emerge in these collective excitonic bands. To see this, we break time-reversal symmetry by adding an out-of-plane external magnetic field $B$, which (together with the magnetic dipole $\mu_{B}$) introduces a Zeeman splitting $\hat{H}\to\hat{H}^{\prime}\equiv\hat{H}+\mu_{B}B\sum_{\tau=\pm,\bm{R}}\tau\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$ between the valley-degenerate doublets, and compute the Berry curvature $\Omega({\bm{Q}})=i\nabla_{\bm{Q}}\times\langle\bm{Q}|\nabla_{\bm{Q}}|\bm{Q}\rangle$ for each band at several $\nu_{e}$, where $|\bm{Q}\rangle$ are right eigenstates of $\hat{H}^{\prime}$  [^13] Here, without loss of generality, we choose the right eigenvectors of the non-Hermitian Hamiltonian to define the Berry curvature [70, 71]. We leave a rigorous study of non-Hermitian topology for future work. We specifically focus on $\Omega({\bm{Q}})$ of (directly) optically accessible states (i.e., within the LC) and compute the numerical results for $\nu_{e}=0$ and $\frac{1}{2}$ in Fig. 4(a), evaluated under a standard momentum discretization scheme [51].
@@
At $\nu_{e}=0$, both collective excitonic states (labeled with respect to their energies by $\Lambda=0,1$) exhibit nontrivial and opposite $\Omega({\bm{Q}})$ centered at $\bm{Q}=0$. Notably, the summation of $\Omega({\bm{Q}})$ in LC, $C_{\mathrm{LC}}$, for these bands takes integer values $\mp 1$. Such a property results from the phase winding of the light-mediated coupling between the valley-doublets, which appears as the off-diagonal terms in the effective low momentum approximation of $\hat{H}^{\prime}$, denoted as $\hat{h}(\bm{Q})$ [31]. With $\bm{Q}=Q(\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y})$, where $\phi$ is the polar angle of $\bm{Q}$, we have:
@@
In contrast, $C_{\mathrm{LC}}$ from generic arrays (see table 1) may be trivial if the low-$Q$ model of the target band is different from Eq. (7). For instance, both collective excitonic bands from the rectangular lattice at $\nu_{e}=\frac{1}{2}$ give zero Berry curvature, as shown in Fig. 4(a), and hence $C_{\mathrm{LC}}=0$. This is due to the absence of $C_{3}$ rotational symmetry, leading to nonzero off-diagonal terms in the corresponding Hamiltonian at $Q=0$, which indicates a gap separating two topologically trivial bands [31].
Figure 4: Comparison between properties of topological and non-topological collective bands from $\nu_{e}=0$ and $\nu_{e}=\frac{1}{2}$, respectively, with Zeeman splitting $\mu_{B}B=20\gamma$. (a) Berry curvature ($\Omega$) distribution in Bloch momentum magnitude $Q$ for the two bands at each $\nu_{e}$ (labeled by colors and $\Lambda$), with solid and dashed lines indicating results from the full model $\hat{H}^{\prime}$ and the low-momentum model Eq. (7), respectively. The vertical axis is in units of $J/\gamma$, where $J$ is fit from $\Gamma_{\bm{Q}}$ in Fig. 2(a) using a momentum sample within $Q\leq{\pi}/{25\lambda_{\mathrm{ex}}}$. The gray vertical line indicates the momentum satisfying $JQ^{2}=\mu_{B}B$. The inset shows that the dependence of $\Omega$ on the polar angle of $\bm{Q}$ is negligible. (b) The phase of $S_{-+}$ at $k=\frac{2\pi}{\lambda_{\mathrm{ex}}}$ and $\bm{k}_{||}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$ for each $\nu_{e}$. All other parameters are the same as in Fig. 2.
@@
Extracting band properties via optical reflectivity.— Finally, we discuss optical experiments to probe the aforementioned collective excitonic properties. More specifically, we illustrate that information about $\hat{H}^{\prime}$ can be obtained by measuring the reflection sector of the far-field scattering matrix for plane-wave incident light (near exciton resonances). For Bravais lattices (e.g., triangular and rectangular arrays), components of this matrix satisfy:
@@
where $\bm{k}_{||}$ is the in-plane component of $\bm{k}$ and $\tau$ ($\tau^{\prime}$) stands for the (circular) polarization of light here (which is locked to the valley index of electrons in TMDs [6]). $g_{\tau,\tau^{\prime}}(k,\bm{Q})=-\frac{i}{2\mathcal{A}_{\nu_{e}}}\frac{k^{2}\delta_{\tau,\tau^{\prime}}-(\bm{e}_{\tau}^{\ast}\cdot\bm{Q})(\bm{e}_{\tau^{\prime}}\cdot\bm{Q})}{k^{2}\sqrt{k^{2}-Q^{2}}}$ is the Fourier transform of the Green’s tensor [31]. $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})=\langle\hat{x}_{\bm{Q},\tau}(\omega-\hat{H}^{\prime})^{-1}\hat{x}_{\bm{Q},\tau^{\prime}}^{\dagger}\rangle$ describes the collective response susceptibility of the exciton, evaluated in the vacuum state, with $\hat{x}_{\bm{Q},\tau}$ denoting the Fourier transform of $\hat{x}_{\bm{R},\tau}$. As shown in Eq. (8), measurements of $S_{\tau,\tau^{\prime}}(\bm{k})$ can help us determine $\hat{H}^{\prime}$ and the corresponding excitonic properties within the LC. In particular, in addition to standard spectral properties such as collective lineshifts and linewidths, we find that $C_{\mathrm{LC}}$ can be extracted from (half of) the winding number of phase of $S_{\tau,-\tau}$ with respect to the polar angle of $\bm{k}_{||}$, as shown in Fig. 4(b).
@@
Outlook. — Our formalism can be further generalized to describe the collective behavior of moiré excitons involving other strong correlations to study their interplay. For instance, going beyond the single excitation subspace, nonlinearities inherent in these emitters can lead to a variety of interesting physical phenomena, including Dicke superradiance [53, 54], optical bistability [55], leaky condensation [56], and phase space filling [57, 58, 59]. Another intriguing problem is whether the spin correlations that emerge near half-filling $\nu_{e}=1$, such as the magnetic polaron effect [60, 61] and kinetic magnetism [62], could play a role in the cooperative excitonic properties.
Acknowledgements. — We acknowledge A. Srivastava, S. Yelin and A. Asenjo-Garcia and D. Goncalves for useful discussions. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. D.E.C. acknowledges support from the European Union, under European Research Council grant agreement No 101002107 (NEWSPIN), FET-Open grant agreement No 899275 (DAALI) and EIC Pathfinder Grant No 101115420 (PANDA); the Government of Spain (Severo Ochoa Grant CEX2019-000910-S [MCIN/AEI/10.13039/501100011033]); QuantERA II project QuSiED, co-funded by the European Union Horizon 2020 research and innovation programme (No 101017733) and the Government of Spain (European Union NextGenerationEU/PRTR PCI2022-132945 funded by MCIN/AEI/10.13039/501100011033); Generalitat de Catalunya (CERCA program and AGAUR Project No. 2021 SGR 01442); Fundació Cellex, and Fundació Mir-Puig.

### v2

@@ Appendix A End Matter
Appendix A: Details of Berry curvature. — Incorporating an out-of-plane external magnetic field $B$ adds a Zeeman splitting between the valley-degenerate doublets to Eq. (4). We accordingly proceed with the Hamiltonian $\hat{H}\to\hat{H}^{\prime}\equiv\hat{H}+\mu_{B}B\sum_{\tau=\pm,\bm{R}}\tau\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$, where $\mu_{B}$ denotes the magnetic dipole. The Berry curvature $\Omega({\bm{Q}})=i\nabla_{\bm{Q}}\times\langle\bm{Q}|\nabla_{\bm{Q}}|\bm{Q}\rangle$ is computed for each band at several $\nu_{e}$, where $|\bm{Q}\rangle$ are right eigenstates of $\hat{H}^{\prime}$. Note that without loss of generality, the right eigenvectors of the non-Hermitian Hamiltonian are chosen to define the Berry curvature [81, 82] (a more rigorous study of non-Hermitian topology is left for future work). We specifically focus on $\Omega({\bm{Q}})$ of (directly) optically accessible states (i.e., within the LC) and compute the numerical results for $\nu_{e}=0$ and $\frac{1}{2}$ [see Fig. 4(a)], evaluated under a standard momentum discretization scheme [83].
@@
At $\nu_{e}=0$, both collective excitonic states (labeled with respect to their energies by $\Lambda=0,1$) exhibit nontrivial and opposite $\Omega({\bm{Q}})$ centered at $\bm{Q}=0$. Notably, the summation of $\Omega({\bm{Q}})$ in LC, $C_{\mathrm{LC}}$, for these bands takes integer values $\mp 1$. As briefly mentioned in the main text, topological features here originates from the phase winding of $\bm{e}_{\tau}^{\ast}\cdot\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)\cdot\bm{e}_{-\tau}$, which appears as the off-diagonal terms in the effective low momentum approximation of $\hat{H}^{\prime}$, denoted as $\hat{h}(\bm{Q})$ [38]. With $\bm{Q}=Q(\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y})$, where $\phi$ is the polar angle of $\bm{Q}$, we have:
@@
In contrast, $C_{\mathrm{LC}}$ from generic arrays (see table 1) may be trivial if the low-$Q$ model of the target band is different from Eq. (7). For instance, both collective excitonic bands from the rectangular lattice at $\nu_{e}=\frac{1}{2}$ give zero Berry curvature, and hence $C_{\mathrm{LC}}=0$. This is due to the absence of $C_{3}$ rotational symmetry, leading to nonzero off-diagonal terms in the corresponding Hamiltonian at $Q=0$, which indicates a gap separating two topologically trivial bands [38].
@@
Figure 4: Comparison between properties of topological and non-topological collective bands from $\nu_{e}=0$ and $\nu_{e}=\frac{1}{2}$, respectively, with Zeeman splitting $\mu_{B}B=20\gamma$. (a) Dependence of Berry curvature ($\Omega$) on Bloch momentum magnitude $Q$ for the two bands at each $\nu_{e}$ (labeled by colors and $\Lambda$), with solid and dashed lines indicating results from the full model $\hat{H}^{\prime}$ and the low-momentum model Eq. (7), respectively. The vertical axis is in units of $J/\gamma$, where $J$ is fit from $\Gamma_{\bm{Q}}$ in Fig. 2(a) using a momentum sample within $Q\leq{\pi}/{25\lambda_{\mathrm{ex}}}$. The gray vertical line indicates the momentum satisfying $JQ^{2}=\mu_{B}B$. The inset shows that the dependence of $\Omega$ on the polar angle of $\bm{Q}$ is negligible. (b) The phase of $S_{-+}$ at $k=\frac{2\pi}{\lambda_{\mathrm{ex}}}$ and $\bm{k}_{||}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$ for each $\nu_{e}$. All other parameters are the same as in Fig. 2.
Appendix B: Extracting band properties via optical reflectivity. — In this appendix, we discuss optical experiments to probe the aforementioned collective excitonic properties. The numerical results of Berry curvature for $\nu_{e}=0$ and $\frac{1}{2}$ are shown in Fig. 4(a), and their manifestation in optical measurements is illustrated in Fig. 4(b). Here, far-field reflection experiment with plane-wave incident light near exciton resonances is considered, and the target physical observables are the amplitude and phase of the scattered field, which can be obtained via standard homodyne technique [29]. Below we illustrate that information about $\hat{H}^{\prime}$ can be inferred from the scattering matrix extracted from such measurement. For Bravais lattices (e.g., triangular and rectangular arrays), components of this matrix satisfy:
@@
where $\bm{k}_{||}$ is the in-plane component of $\bm{k}$ and $\tau$ ($\tau^{\prime}$) stands for the (circular) polarization of light here (which is locked to the valley index of electrons in TMDs [6]). $g_{\tau,\tau^{\prime}}(k,\bm{Q})=-\frac{i}{2\mathcal{A}_{\nu_{e}}}\frac{k^{2}\delta_{\tau,\tau^{\prime}}-(\bm{e}_{\tau}^{\ast}\cdot\bm{Q})(\bm{e}_{\tau^{\prime}}\cdot\bm{Q})}{k^{2}\sqrt{k^{2}-Q^{2}}}$ is the Fourier transform of the Green’s tensor [38]. $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})=\langle\hat{x}_{\bm{Q},\tau}(\omega-\hat{H}^{\prime})^{-1}\hat{x}_{\bm{Q},\tau^{\prime}}^{\dagger}\rangle$ describes the collective response susceptibility of the exciton, evaluated in the vacuum state, with $\hat{x}_{\bm{Q},\tau}$ denoting the Fourier transform of $\hat{x}_{\bm{R},\tau}$. Once $\hat{H}^{\prime}$ is reconstructed from $S_{\tau,\tau^{\prime}}(\bm{k})$ via Eq. (8), the corresponding spectral and topological excitonic properties within the LC could be determined. In particular, $C_{\mathrm{LC}}$ can be extracted from (half of) the winding number of phase of $S_{\tau,-\tau}$ with respect to the polar angle of $\bm{k}_{||}$, as shown in Fig. 4(b).